OpenStudy (kittygirl99):

x+7/8=4 what does x means

OpenStudy (kittygirl99):

help

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

x=25/8

OpenStudy (anonymous):

@kittygirl99

OpenStudy (kittygirl99):

kk

OpenStudy (anonymous):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : x+7/8-(4)=0 7 Simplify — 8 Rewrite the whole as a fraction using 8 as the denominator : x x • 8 x = — = ————— 1 8 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: x • 8 + 7 8x + 7 ————————— = —————— 8 8 Rewrite the whole as a fraction using 8 as the denominator : 4 4 • 8 4 = — = ————— 1 8 Adding up the two equivalent fractions (8x+7) - (4 • 8) 8x - 25 ———————————————— = ——————— 8 8 Where a fraction equals zero, its numerator, the part which is above the fraction line, must equal zero. Now,to get rid of the denominator, I would multiply both sides of the equation by the denominator. Here's how: 8x-25 ————— • 8 = 0 • 8 8 Now, on the left hand side, the 8 cancels out the denominator, while, on the right hand side, zero times anything is still zero. The equation now takes the shape : 8x-25 = 0 Solve : 8x-25 = 0 Add 25 to both sides of the equation : 8x = 25 Divide both sides of the equation by 8: x = 25/8 = 3.125 Your answer would now be x = 25/8 = 3.125.